OpenStudy (anonymous):

If the sequence is geometric, what is the common ratio? 6, 18, 54, 162, ... 12 3 27 not geometric

OpenStudy (cwrw238):

18/6 = 3

OpenStudy (anonymous):

\[ \Huge 3 \]

OpenStudy (anonymous):

clearly 3

OpenStudy (hoblos):

18/6 = 54/18 = 162/54 = 3

OpenStudy (anonymous):

3

OpenStudy (hoblos):

so it is geometric of ratio 3

OpenStudy (anonymous):

common ratio.